# UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers

### v1

@@ Abstract
Neural NLP models are often miscalibrated, assigning high confidence to incorrect predictions, which undermines selective prediction and high-stakes deployment. Post-hoc calibration methods adjust output probabilities but leave internal computation unchanged, while ensemble and Bayesian approaches improve uncertainty at substantial training or storage cost. We propose UAT-Lite, an inference-time framework that makes self-attention uncertainty-aware using approximate Bayesian inference via Monte Carlo dropout in pretrained transformer classifiers. Token-level epistemic uncertainty is estimated from stochastic forward passes and used to modulate self-attention during contextualization, without modifying pretrained weights or training objectives. We additionally introduce a layer-wise variance decomposition to diagnose how predictive uncertainty accumulates across transformer depth. Across the SQuAD 2.0 answerability, MNLI, and SST-2, UAT-Lite reduces Expected Calibration Error by approximately 20% on average relative to a fine-tuned BERT-base baseline while preserving task accuracy, and improves selective prediction and robustness under distribution shift. 111github.com/eliashossain001/uq_decomposition
UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers
@@ 1 Introduction
Pretrained transformer-based language models have achieved state-of-the-art performance across a wide range of NLP tasks, from question answering to clinical decision support Devlin et al. (2019); Brown et al. (2020). Despite these advances, modern transformers exhibit systematic miscalibration, frequently assigning high confidence to incorrect predictions Desai and Durrett (2020). This misalignment between confidence and correctness is particularly problematic for selective prediction and related high-stakes applications.
@@
Fig.˜1 summarizes existing approaches to uncertainty estimation. Standard transformer inference is deterministic and does not represent uncertainty during internal computation (Fig.˜1A). Post-hoc calibration methods, such as temperature scaling, adjust confidence only after representations are formed, leaving internal reasoning unchanged (Fig.˜1B; Guo et al., 2017). Stochastic approaches, including Monte Carlo (MC) dropout and deep ensembles (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017), improve uncertainty estimates but largely treat uncertainty as an output-level signal, without influencing attention.
@@
This limitation reflects a broader gap: transformers are trained to optimize predictive accuracy rather than confidence reliability Guo et al. (2017). While post-hoc methods are lightweight and effective in-domain, they do not alter token interactions. Bayesian approaches integrate uncertainty more deeply but typically require architectural modification or specialized training, limiting compatibility with pretrained models. This raises the following question: can epistemic uncertainty shape a transformer’s internal attention at inference time, without retraining or modifying pretrained weights?
@@
We argue that effective uncertainty quantification requires uncertainty to actively shape attention, rather than merely annotate final predictions. For ambiguous inputs, a calibrated model should hedge internal reasoning instead of propagating early attention decisions deterministically. Existing stochastic inference methods expose uncertainty only at the output level and do not use it to modulate internal attention patterns (Fig.˜1C).
@@
Motivated by this observation, we propose UAT-Lite, an inference-time framework for encoder-based pretrained transformer classifiers. UAT-Lite estimates token-level epistemic uncertainty via Monte Carlo sampling and uses it to downweight attention contributions from unstable tokens, without modifying pretrained weights or training objectives. Our main contributions are:
@@
Uncertainty-Weighted Attention. An inference-time mechanism that modulates self-attention using token-level epistemic uncertainty estimated via Monte Carlo sampling.
@@
Layer-Wise Uncertainty Attribution. A diagnostic variance decomposition that attributes predictive uncertainty across transformer depth.
Comprehensive Evaluation. Empirical evaluation across multiple NLP benchmarks demonstrating improved calibration, selective prediction, and robustness under distribution shift without degrading task performance.
UAT-Lite introduces no additional trainable parameters and operates entirely at inference time, requiring only a small number of stochastic forward passes. Post-hoc calibration methods such as temperature scaling are complementary rather than competing. The framework trades modest inference-time cost for improved reliability while avoiding the training and storage overhead of ensemble-based approaches. Our study focuses on encoder-based transformers; extending uncertainty-aware attention to decoder-only generative models is left for future work.
@@ 2 Preliminaries
Bayesian inference and uncertainty.
Bayesian analysis incorporates parameter uncertainty by averaging predictions over probable parameter values inferred from the training data. This perspective distinguishes epistemic uncertainty, which arises from limited data or insufficient coverage and reflects uncertainty about the model itself, from aleatoric uncertainty, which represents irreducible noise in the observations. In large neural models, epistemic uncertainty becomes particularly significant in the presence of semantic ambiguity or distributional shifts, where confident yet incorrect predictions frequently occur.
Monte Carlo (MC) Dropout.
We briefly recall MC Dropout as an approximate Bayesian inference technique used throughout this work. MC dropout retains dropout layers at inference time and performs multiple stochastic forward passes, yielding a predictive distribution whose variance estimates epistemic uncertainty and whose mean defines the final prediction. Conventional MC dropout treats uncertainty solely at the output level and does not influence internal representations or attention mechanisms, motivating methods that integrate uncertainty directly into the inference process.
Token-level uncertainty as a proxy signal.
While epistemic uncertainty is often measured at the output level, internal transformer computation is structured around token-level representations and attention interactions. Consequently, uncertainty estimates derived from stochastic embedding representations can serve as a lightweight proxy for identifying unstable or ambiguous token contributions. This perspective motivates the use of token-level uncertainty as a control signal for modulating attention during inference, without requiring architectural changes or retraining.
3 Related Work
Uncertainty estimation and calibration for neural networks have been widely studied, with post-hoc methods such as temperature scaling and Platt scaling (Guo et al., 2017; Platt et al., 1999) providing simple and effective confidence correction without altering model internals. Ensemble-based approaches, including Deep Ensembles (Lakshminarayanan et al., 2017), yield strong calibration but incur substantial training and storage overhead. Bayesian neural networks, stochastic inference methods, and recent Bayesian Transformer variants (Blundell et al., 2015; Gal and Ghahramani, 2016; Ritter et al., 2018; Fan et al., 2020; Zhang et al., 2021; Chen and Li, 2023) offer principled uncertainty modeling but typically require architectural modification, variational training, or task-specific retraining. In contrast, UAT-Lite integrates epistemic uncertainty directly into self-attention at inference time, preserving pretrained weights and standard training pipelines while enabling uncertainty-aware internal computation. A comprehensive discussion of related work and detailed comparisons is provided in App.˜G.
@@ 4 UAT-Lite
UAT-Lite is an uncertainty-aware extension of pretrained transformer classifiers designed to improve confidence calibration at inference time without modifying model architecture, training objectives, or learned parameters. The framework integrates approximate Bayesian uncertainty estimation into internal attention computation, allowing epistemic uncertainty to actively influence contextualization while preserving the base transformer’s parameterization. In addition, UAT-Lite introduces a diagnostic layer-wise variance decomposition that characterizes how predictive uncertainty evolves across transformer depth, enabling post-hoc analysis without affecting the forward pass. Final predictions are produced via stochastic inference with confidence-aware decision shaping. Overall, UAT-Lite preserves the training cost and parameter count of the underlying encoder, incurring only modest inference-time overhead.
@@ 4.1 Problem Setup and Notation
We consider an encoder-based pretrained transformer for classification. Given an input sequence , the model maps through stacked layers to contextual representations , which are converted to a predictive distribution by a task-specific classification head. In standard evaluation mode, pretrained transformers produce a single point estimate: dropout is disabled and the same input yields the same representations and prediction given fixed parameters. While effective for accuracy, this design provides no mechanism for epistemic uncertainty. Our goal is to introduce controlled inference-time stochasticity that allows uncertainty to influence internal attention without modifying training or parameters.
@@ 4.2 Estimating Uncertainty via MC Sampling
We estimate predictive uncertainty using MC dropout, a lightweight approximation to Bayesian inference. Dropout layers are retained during inference and stochastic forward passes are performed, inducing variability via random neuron masking. Aggregating these samples yields a predictive mean and an estimate of epistemic variability across stochastic passes during evaluation.
@@
Following Kendall and Gal (2017), predictive uncertainty decomposes into aleatoric and epistemic components. MC dropout primarily captures epistemic uncertainty through approximate posterior sampling during inference.
@@
To balance representational stability and uncertainty estimation, we apply component-specific dropout during inference (0.1 for embeddings, 0.2 for attention layers, and 0.3 for feed-forward layers). These rates follow common practice in transformer fine-tuning and are kept fixed across tasks; calibration trends are stable under moderate variations of these values.
@@
We use a single-pass MC inference procedure in which the same stochastic forward passes are used both to (i) estimate token-level embedding uncertainty and (ii) obtain the final MC predictive distribution. As detailed in Algorithm 1, for each pass we sample dropout masks, compute embeddings , propagate the input through the encoder with uncertainty-weighted attention, and record logits for aggregation. Token-level uncertainty is computed by aggregating embedding variability across these passes and is reused consistently within the same inference run. This design avoids any separate embedding-only sampling stage and ensures that inference-time overhead scales linearly with .
@@ 4.3 Uncertainty-Weighted Self-Attention
We incorporate epistemic uncertainty directly into self-attention by modulating attention logits using token-level uncertainty estimates computed at the embedding layer during contextualization.
@@
For each input token , we obtain stochastic embedding samples . Token-level uncertainty is defined as
where denotes the embedding dimensionality. We interpret as an embedding-level proxy for epistemic uncertainty induced by dropout.
This quantity is computed once per input and shared across all layers and attention heads. The computational cost is negligible relative to Monte Carlo methods, as it involves only per-token operations over the sampled embeddings and does not require additional forward passes.
Scale sensitivity.
The magnitude of remains well-behaved due to embedding normalization and averaging across dimensions. Sensitivity analysis over and (§˜D.1) shows stable calibration across a broad range of values; explicit normalization is therefore unnecessary.
@@
Let , , and denote query, key, and value projections. Standard scaled dot-product attention logits are
We modulate these logits using token-level uncertainty:
where is a fixed uncertainty penalty. Attention probabilities are obtained via softmax:
@@
Design choices.
Uncertainty modulation is applied uniformly across layers and heads and affects only the key dimension in the attention mechanism. Unless otherwise stated, the uncertainty penalty (default ) is set once on development data and held fixed across tasks. §˜D.1 shows that calibration is stable across a broad range of and . The method introduces no additional trainable parameters beyond the base model.
@@ 4.4 Layer-Wise Uncertainty Attribution
We analyze how predictive uncertainty propagates across transformer depth under inference-time stochasticity. This analysis is diagnostic and does not affect forward computation or model behavior.
@@
Using the law of total variance, total predictive variance can be decomposed into layer-specific contributions under inference-time stochasticity. A formal derivation is provided in App.˜A.
Theorem 1 (Layer-Wise Variance Decomposition).
Under inference-time stochasticity, total predictive variance can be expressed as the sum of layer-specific variance contributions.
@@
This result follows from standard variance identities and is not a novel mathematical contribution. Its value lies in enabling diagnostic inspection of uncertainty evolution across transformer depth, which is unavailable in post-hoc calibration methods. Layer-wise contributions should be interpreted as approximate signals rather than strictly causal attributions.
@@ 4.5 Confidence-Aware Decision Shaping
MC dropout produces a distribution of stochastic predictions. The predictive mean is computed from the stochastic forward passes. Post-hoc temperature scaling is applied to the Monte Carlo mean logits, with the temperature parameter learned on the validation set to calibrate predictive confidence.
@@
Uncertainty estimates enable abstention via confidence thresholds selected on the validation split of each dataset and applied uniformly at test time across all evaluated methods.
@@ 5.1 Datasets
We evaluate on five benchmarks spanning general NLP and clinical domains commonly used in prior calibration and robustness studies. All experiments focus exclusively on encoder-based pretrained transformer classifiers (BERT-family backbones), rather than decoder-only language models. General NLP datasets include SQuAD 2.0 Rajpurkar et al. (2018) (QA with unanswerable questions), MNLI Williams et al. (2018) (natural language inference with matched/mismatched splits for distribution shift analysis), and SST-2 Socher et al. (2013) (binary sentiment classification). Clinical benchmarks include MedQA Jin et al. (2021) and PubMedQA Jin et al. (2019), which are used to assess domain transfer rather than clinical validity.

### v2

@@ Abstract
Neural NLP models are often miscalibrated and overconfident, assigning high confidence to incorrect predictions and failing to express uncertainty during internal evidence aggregation. This undermines selective prediction and high-stakes deployment. Post-hoc calibration methods adjust output probabilities but leave internal computation unchanged, while ensemble and Bayesian approaches improve uncertainty at substantial training or storage cost. We propose UAT-Lite, an inference-time framework that makes self-attention uncertainty-aware via Monte Carlo dropout in pretrained transformer classifiers. Unlike output-level calibration (e.g., TS), UAT-Lite injects epistemic uncertainty directly into attention, enabling uncertainty-aware routing during contextualization and token-level diagnostic signals beyond global logit rescaling. Token-level epistemic uncertainty is estimated from stochastic forward passes and used to modulate self-attention during contextualization, without modifying pretrained weights or training objectives. We additionally introduce a layer-wise variance decomposition to diagnose how predictive uncertainty accumulates across transformer depth. Across SQuAD 2.0 answerability, MNLI, and SST-2, UAT-Lite achieves an average relative ECE reduction of approximately 20% compared with a fine-tuned BERT-base baseline while preserving accuracy, and yields more informative uncertainty behavior for selective prediction under distribution shift.
111github.com/eliashossain001/uq_decompositionUAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers
@@ 1 Introduction
Pretrained transformer-based language models achieve state-of-the-art performance across a wide range of NLP tasks, from question answering to clinical decision support Devlin et al. (2019); Brown et al. (2020). Despite these advances, modern transformers exhibit systematic miscalibration, often assigning high confidence to incorrect predictions Desai and Durrett (2020). This mismatch between confidence and correctness is particularly problematic for selective prediction and other high-stakes settings.
@@
Fig.˜1 summarizes common approaches to uncertainty estimation. Standard transformer inference is deterministic and does not represent uncertainty in internal computation (Fig.˜1A). Post-hoc calibration methods, such as temperature scaling (TS), adjust confidence only after representations are formed, leaving internal computation unchanged (Fig.˜1B; Guo et al., 2017). Stochastic approaches, including Monte Carlo (MC) dropout and deep ensembles (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017), improve uncertainty estimates but typically treat uncertainty as an output-level signal without influencing attention.
@@
This gap reflects a broader issue: transformers are trained to optimize predictive accuracy rather than confidence reliability Guo et al. (2017). While post-hoc methods are lightweight and effective in-domain, they do not alter token interactions. Bayesian approaches integrate uncertainty more deeply but often require architectural changes or specialized training, which limits compatibility with pretrained models. This motivates a central question: can epistemic uncertainty shape a transformer’s attention at inference time without retraining or modifying pretrained weights?
@@
We argue that effective uncertainty quantification should allow uncertainty to shape attention, not merely annotate final predictions. For ambiguous inputs, a reliable model should modulate evidence aggregation rather than propagate early attention decisions deterministically. Existing stochastic inference methods expose uncertainty mainly at the output level and do not use it to modulate internal attention patterns (Fig.˜1C). Although temperature scaling is highly effective for output calibration, it only rescales logits and cannot provide token-level epistemic signals or uncertainty-aware routing inside the model.
@@
Motivated by these observations, we propose UAT-Lite, an inference-time framework for encoder-based pretrained transformer classifiers. UAT-Lite estimates token-level epistemic uncertainty via MC dropout and uses it to downweight attention contributions from unstable tokens during contextualization, without modifying pretrained weights or training objectives. Our main contributions are as follows:
@@
Uncertainty-Weighted Attention. An inference-time mechanism that injects token-level epistemic uncertainty, estimated via MC dropout, into self-attention, downweighting unstable token contributions during contextualization.
@@
Layer-Wise Uncertainty Attribution. A variance decomposition that attributes predictive uncertainty across transformer depth, enabling diagnostic analysis of where uncertainty amplifies during inference.
Comprehensive Evaluation. Experiments on SQuAD 2.0 answerability, MNLI, and SST-2 that show improved calibration and selective prediction, along with robustness analyses under distribution shift.
UAT-Lite introduces no additional trainable parameters and operates strictly at inference time, using a controllable Monte Carlo budget (e.g., ) to trade compute for reliability. UAT-Lite is not intended to replace lightweight output-level calibrators such as TS, which is optimized on development logits and often performs best in-domain. While TS rescales final logits after representations are formed, UAT-Lite modulates internal evidence aggregation by making attention uncertainty-aware. Concretely, UAT-Lite provides (i) token-level epistemic signals, (ii) uncertainty-aware routing via attention modulation, and (iii) layer-wise diagnostic attribution; TS provides only output probability rescaling. Because the two mechanisms act at different stages, they are complementary and can be stacked (UAT-Lite +TS) when calibrated probabilities are needed. Overall, the framework trades a controllable inference-time cost for improved reliability while avoiding the training and storage overhead of ensemble-based alternatives. Our study focuses on encoder-based transformers; extending uncertainty-aware attention to decoder-only generative models is left for future work.
2 Related Work
Uncertainty estimation and calibration for neural networks have been studied extensively. Post-hoc methods such as temperature scaling and Platt scaling (Guo et al., 2017; Platt et al., 1999) provide simple and effective confidence correction without altering model internals. Ensemble-based approaches, including Deep Ensembles (Lakshminarayanan et al., 2017), often yield strong calibration performance but incur substantial training and storage overhead. Bayesian neural networks, stochastic inference methods, and recent Bayesian transformer variants (Blundell et al., 2015; Gal and Ghahramani, 2016; Ritter et al., 2018; Fan et al., 2020; Zhang et al., 2021; Chen and Li, 2023) offer principled approaches to uncertainty modeling, but they typically require architectural modifications, variational training, or task-specific retraining. By contrast, UAT-Lite integrates epistemic uncertainty directly into self-attention at inference time, preserving pretrained weights and standard training pipelines while enabling uncertainty-aware internal computation. A more comprehensive discussion of related work and detailed comparisons is provided in App.˜G.
@@ 3 Preliminaries
We briefly review the uncertainty concepts and stochastic inference protocol used throughout the paper. Bayesian perspectives distinguish epistemic uncertainty, which arises from limited data or distribution shift, from aleatoric uncertainty, which reflects irreducible observation noise. In our setting, epistemic effects dominate under semantic ambiguity and distribution shift, where transformers can be confidently wrong. We use MC dropout as a lightweight approximate Bayesian inference method. Specifically, dropout is retained at test time, and we run stochastic forward passes to obtain a predictive distribution whose mean is used for prediction and whose variability serves as an estimate of epistemic uncertainty. Standard MC dropout treats uncertainty as an output-level signal and does not influence internal computation. Because transformer inference is organized around token representations and attention interactions, we use token-level uncertainty derived from stochastic embeddings as a proxy for unstable token contributions. This motivates modulating attention with token uncertainty at inference time, without retraining or changing pretrained weights.
@@ 4 UAT-Lite
UAT-Lite is an inference-time extension of pretrained transformer classifiers that injects epistemic uncertainty into self-attention to improve confidence reliability without modifying the model architecture, training objectives, or learned parameters. The framework uses MC dropout to estimate token-level variability and to modulate attention during contextualization. UAT-Lite also provides a diagnostic layer-wise variance decomposition that characterizes how predictive uncertainty accumulates across transformer depth without affecting the forward pass. Final predictions are produced by aggregating stochastic logits, and optional TS can be applied to the MC-mean logits when calibrated probabilities are required.
@@ 4.1 Problem Setup and Notation
We consider an encoder-based pretrained transformer for classification. Given an input sequence , the model maps through layers to contextual representations , which are converted to a predictive distribution by a task-specific head. In standard evaluation, inference is deterministic: dropout is disabled, and the same input yields the same prediction for fixed parameters. While effective for accuracy, this setting provides no mechanism for representing epistemic uncertainty during internal computation. Our goal is to introduce controlled inference-time stochasticity and use it to modulate attention without retraining or parameter updates.
@@ 4.2 Estimating Uncertainty via MC Sampling
We estimate predictive uncertainty using MC dropout, a lightweight approximation to Bayesian inference. Dropout is retained at test time, and we perform stochastic forward passes, each induced by a different dropout mask. Aggregating the resulting logits yields a predictive mean and an estimate of epistemic variability across passes.
@@
Following Kendall and Gal (2017), predictive uncertainty decomposes into aleatoric and epistemic components. MC dropout primarily captures epistemic uncertainty through approximate posterior sampling at inference time.
@@
To balance variability and stability, we use component-specific dropout rates during inference: 0.1 for embeddings, 0.2 for attention layers, and 0.3 for feed-forward layers. These rates are fixed across tasks, and calibration trends remain stable under moderate rate variations.
@@
We use a single-pass MC procedure in which the same stochastic forward passes are used to (i) estimate token-level embedding uncertainty and (ii) compute the final MC predictive distribution. Concretely, during pass we sample dropout masks and compute embeddings , then update running uncertainty estimates from the embedding samples observed so far. To avoid circular dependence within a pass, we modulate attention in pass using the lagged estimate , with , as summarized in Algorithm 1. After passes, the final token uncertainty is , and predictions are obtained by aggregating logits . This design avoids a separate embedding-only stage and ensures that overhead scales linearly in .
@@ 4.3 Uncertainty-Weighted Self-Attention
We incorporate epistemic uncertainty into self-attention by modulating attention logits using token-level uncertainty estimated at the embedding layer.
@@
For each token , we obtain stochastic embedding samples via MC dropout. We define token uncertainty as
where is the embedding dimension. We treat as an embedding-level proxy for the epistemic uncertainty induced by dropout and use it as a control signal for uncertainty-aware attention routing.
Single-pass / online estimation. To avoid a separate embedding-only sampling stage, we estimate online during the same stochastic passes used for prediction. Specifically, after completing pass , we maintain a running estimate (e.g., via Welford updates), and we use to modulate attention in pass . After passes, the final estimate aggregates variability across all samples. During the MC run, each pass uses the lagged estimate consistently across all layers and heads.
Scale/geometry diagnostics. We test whether is a trivial artifact of embedding scale or geometry through embedding-rescaling invariance, correlations with output-space epistemic measures (entropy/MI), and uncertainty-definition ablations (§˜D.8).
@@
Let , , and denote the query, key, and value projections. Standard scaled dot-product attention logits are
We define uncertainty-aware logits by attenuating the pre-softmax scores:
where is a fixed penalty parameter and specifies the projection variant.
Positioning vs. TS. UAT-Lite does not replace TS: TS rescales final logits post hoc, whereas UAT-Lite modulates pre-softmax attention logits using token-level epistemic uncertainty, enabling uncertainty-aware routing and token-level diagnostics beyond output rescaling. We analyze sensitivity to and compare penalty forms (exponential vs. linear vs. reciprocal) in §˜D.6.
Projection variants (Q/K/V).
Using , we instantiate as follows:
We then obtain attention probabilities via the softmax:
We additionally consider V-only and QKV by applying the same factor to the value path, , and using as the attention output.
@@
Prior uncertainty-aware attention methods typically require architectural modification or variational training (Chen and Li, 2023; Zhang et al., 2021), whereas our approach operates entirely at inference time. Unless otherwise stated, we default to Q-only, which empirically provides a favorable calibration–accuracy trade-off in our experiments in Section 6.5.
@@ 4.4 Layer-Wise Uncertainty Attribution
We analyze how predictive uncertainty propagates across transformer depth under inference-time stochasticity. This analysis is purely diagnostic and does not alter model behavior.
@@
Using the law of total variance, we decompose the predictive variance under MC dropout into layer-indexed contributions (App.˜A).
Proposition 1 (Diagnostic Layer-Wise Variance Decomposition).
Under the stochastic inference procedure used in this work, predictive variance admits a recursive decomposition across transformer depth via repeated application of the law of total variance.
@@
This decomposition follows from standard variance identities and is not claimed to be a novel theoretical result. Its purpose is to diagnose where uncertainty concentrates or amplifies across depth. The resulting contributions should therefore be interpreted as approximate diagnostic signals rather than causal attributions.
@@ 4.5 Confidence-Aware Decision Shaping
MC dropout yields logits from stochastic passes. We compute the MC-mean logits as and derive the predictive distribution from . When used, post-hoc temperature scaling is fit on the validation set and applied to at test time.
@@
We perform abstention by thresholding confidence on the validation split and applying the same thresholds at test time for all methods.
@@ 5.1 Datasets
We evaluate on five benchmarks spanning general-domain NLP and clinical-language settings that are commonly used in prior calibration and robustness studies. All experiments focus exclusively on encoder-based pretrained transformer classifiers (BERT-family backbones), rather than decoder-only language models. The general-domain datasets include SQuAD 2.0 Rajpurkar et al. (2018), a question-answering benchmark with unanswerable questions; MNLI Williams et al. (2018), a natural language inference benchmark with matched and mismatched splits that support distribution-shift analysis; and SST-2 Socher et al. (2013), a binary sentiment classification benchmark. The clinical-domain benchmarks include MedQA Jin et al. (2021) and PubMedQA Jin et al. (2019), which we use to assess domain transfer rather than clinical validity.